OpenStudy (anonymous):

@ganeshie8 This Is My Last ? On The Test Its Like Derivatives Anyways I Know How To Get Answer I Just Cant Standard Them PLEASE Help

OpenStudy (anonymous):

Derive the equation of the parabola with a focus at (−2, 4) and a directrix of y = 6. Put the equation in standard form. My Answer Is F(x)=1 over 4 (x+2)^2+5 How Do I Put In Standard Form

OpenStudy (anonymous):

\[f(x)= \frac{ 1 }{ 4 }(x+2)^2+5\]

OpenStudy (anonymous):

@ikram002p

OpenStudy (anonymous):

f(x) = one fourthx2 − x + 4 f(x) = −one fourthx2 − x + 4 f(x) = one fourthx2 − x + 5 f(x) = −one fourthx2 − x + 5 These Are The Answer Choices

OpenStudy (anonymous):

HELP?

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

Can Someone HELP

OpenStudy (anonymous):

@zepdrix @Compassionate @tkhunny @aaronq

OpenStudy (anonymous):



OpenStudy (anonymous):

THANK YOU SOO MUCH!

OpenStudy (tkhunny):

Listen to surjithayer. It's called "The Definition". With this information, you NEVER need to worry about having nowhere to go.